OpenStudy (anonymous):

Solve: 4(2x − 1) − 3 = 9(x − 4)

OpenStudy (anonymous):

8x-4-4=9x-36 8x-8=9x-36 -x=-28

OpenStudy (anonymous):

x=28

OpenStudy (anonymous):

x doesn't = 28, that's not an option either. lol

OpenStudy (anonymous):

8x-4-3 -9x +36=0 -x = -36+4+3 -x = -29 x=29

OpenStudy (aroub):

x=29